Wednesday, October 22, 2014

Night Log 2014-10-21

SA: Brent
SO: Veronica
Others: Ockert, Jonathan, Luke (remotely)

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Some short on sky tests at start of night to ensure telescope and HRS are back online. Will resume normal science operations tomorrow night. Data taken for:

2014-1-UNC-001 P1 HRS (Emission line kinematics of RESOLVE galaxies for tracing dark matter)
2014-1-AMNH-005 P4 SCAM (Studying the long-term behavior of classical novae after their eruptions)

Wednesday, October 15, 2014

Night Log 2014-10-14

SA: David
SO: Fred
Others: Hitesh, Deon

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


SAMS Engineering tests

Tuesday, October 14, 2014

Night Log 2014-10-13

SA: David
SO: Fred
Others: Hitesh, Deon

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Engineering tests:
pointing
SAMS

Monday, October 13, 2014

Night Log 2014-10-12

SA: Encarni
SO: Thea
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Telescope offline for engineering week. No observations tonight.

Sunday, October 12, 2014

Night Log 2014-10-11

SA: Encarni
SO: Thea
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Telescope offline due to engineering week. No observations tonight.

Wednesday, October 8, 2014

Night Log 2014-10-07

SA: Alexei
SO: Fred
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


=> Tracker is off tonight. No observations.

Tuesday, October 7, 2014

Night Log 2014-10-06

SA: Alexei
SO: Fred
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


=> Stable night with moderate seeing and low humidity.

=> Have some technical issues.

=> Data were taken for the following programs:
2014-1-DDT-003 P0 HRS (Study of a helium binary in outburst. )
2014-1-POL_RSA_AMNH-001 P1 HRS (Looking for evidence of massive white dwarfs in symbiotic novae)
2014-1-HET_UC-001 P1 HRS (Searching for the Earth-mass planet around alpha Cen B)
2014-1-RSA_HET-002 P2 HRS (Investigating the birth and evolution of X-ray binaries through their radial motion.)
2014-1-DC-001 P2 HRS (Gravitational redshifts in planetary nebulae and white dwarfs and spectra of dwarf galaxies)
2014-1-AMNH-005 P4 SCAM (x3) (Studying the long-term behavior of classical novae after their eruptions)